As studied, the average number of hours spent by senior high school students for their online classes a week is 25 hours with a standard deviation of 4 hours. Assuming that the study is true and the data is normally distributed. Find the probability that a randomly selected senior high school srudent spends less than 21 hours or greater than 30 hours.


1
Expert's answer
2021-05-17T16:01:40-0400

Let "X=" the number of hours spent by senior high school students for their online classes a week: "X\sim N(\mu, \sigma^2)"

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

Given "\mu=25\ h, \sigma=4\ h."


"P((X<21)\cup(X>30))"

"=P(X<21)+1-P(X\leq30)"

"=1+P(Z<\dfrac{21-25}{4})-P(Z\leq\dfrac{30-25}{4})"

"=1+P(Z<-1)-P(Z\leq1.25)"

"\approx1+0.158655-0.894350\approx0.2643"

The probability that a randomly selected senior high school srudent spends less than 21 hours or greater than 30 hours is 0.2643.